Grade 12th passAlgebra

If 14 persons are seated at a round table, then the number of
ways of selecting two persons out of them who are not
seated adjacent to each other is

vamsi
12 Years agoGrade 12th pass

2 Answers

India VK
12 Years ago

First person can be chosen in 14 ways. After the first person is chosen, the second person cannot be chosen from a place adjacent to him. So two places out of remaining 13 cannot be selected. So 11 is the number of ways of selecting the second person.

 

Total number of ways = 14 * 11 = 154

vamsi
12 Years ago
First person can be chosen in 14 ways. After the first person is chosen, the second person cannot be chosen from a place adjacent to him. So two places out of remaining 13 cannot be selected. So 11 is the number of ways of selecting the second person. Total number of ways = 14 * 11 = 154